Mitch made two 10-inch pizzas using a recipe that calls for 3 cups
of flour and 3/4 cup of water. Mitch wants to make more pizzas
without changing the ratio of flour to water.
Complete the sentence about the ingredients Mitch needs to make
eight 10-inch pizzas. Enter a number that makes the sentence
true.
To make eight 10-inch pizzas, Mitch uses 12 cups of flour and
__?__ cups of water.

Answer by math_tutor2020(3816)


Answer: 3 cups of water

Explanation

To make 2 pizzas, he needs 3 cups of flour.
To make 8 pizzas, he needs 12 cups of flour.
This is a jump of "times 4".

Quadrupling the original amount of water has us go from 3/4 cups to (3/4)*4 = 3 cups

Another way to find this answer is to solve this proportion
3/0.75 = 12/x